# DreamPBR: Text-driven Generation of High-resolution SVBRDF with Multi-modal Guidance

### v1

@@ 4.2. Generation Results
DreamPBR is capable of generating realistic or magic materials with only descirptions. In order to demonstrates the abitlity to synthesis wide materials, we obtain a mount of descriptions of materials in by LLM for each type, which are used to sample materials with DreamPBR. The generated textures are enhanced by the super-resolution module and are then rendered as shown in Figure 3. In our sampled 400 textures, they show high consistency with text and the mean of CLIP Score between rendering images and given prompts is 30.198. Besides the consistency of text and images, the diversity of results is quite important for text-driven generative models as well. As demonstrated in Figure 7, we further sample several textures with the same prompt but different random seeds, DreamPBR succeeds in producing diverse textures that follow the descriptions we specify.
@@ 4.2.1. Tileable texture generation
Although the users would introduce various controls, we can generate seamless tileable textures all the time, which allows users to apply the generated textures in different sacles and different scenes. In Figure 8, we presents several tileable textures from direct and guided generation with their splicing results, showing that the effectiveness of circular padding in our method as mentioned in Section 3.2.
@@ 4.2.2. Results of Pixel Control
By finetuning an additional ControlNet, DreamPBR is able to generate textures according to given patterns. In practice, a designer could decide on a pattern in advance, and then try different materials. It may also be the other way around. For those two situations, DreamPBR ensures reasonable texutres for certain patterns or materials as demonstrated in Figure 4 and Figure 5.
With addtional control of binary images, inpainting is also an usual method for users to obtain specified results so that we present several inpainting results in Figure 9 to replace a regions in texture with another regions users descirpt.
@@ 4.2.3. Results of Style Control
A styled image express more easily for a person than only text like Su et al. (2021) does. To do so, we evaluate the adaptation of Su et al. (2021) for our Style Control. Specifically, we obtain several styled images online and present the generation results under different styles from images as shown in Figure 6. Figure 10 illustrates the situation that users would like to combine Style Control with Pixel Control, which enable users to generate the results they want more freely.
@@ 4.2.4. Results of Shape Control
With the ability to generate various textures, DreamPBR can be extended to non-planar objects such as chairs. Specifically by giving a segmented objects, we can utilize dialogue with a large language model to get different descriptions of each regions. For more specified objects, a more direct way is to be in conjunction with cropped areas from exemplar images used with pixel control and style control. Thanks to the tileable features, the results from our pipeline of Shape Control is shown in Figure 11.
@@ 4.3. Comparative Experiments
Leveraging the state of the art generative model, StableDiffusion, DreamPBR is very competitive with previous methods for materials generation. We compare the results generated from DreamPBR of different materials against MaterialGAN  (Guo et al., 2020) and TileGen  (Zhou et al., 2022) in Figure 12. Notably, there are only two categories provided in the compteting methods so that our results are generated by giving prompts, “a PBR material of ground, stone” and “a PBR material of metal”. The comparison shows that DreamPBR can generate textures following distribution of realistic data from dataset like GAN-based methods as well as magic textures from prior information for 2D images.
Moreover, we compare our Pixel Control with those of TileGen in generation with sketches guidance. The comparison results are shown in Figure 13, in which we demonstrate different generation results of TileGen and ours with the same binary masks. DreamPBR surpasses TileGen in sketches-driven generation and shows less artifacts and more precise controls than previous researches on material generation like TileGen.
@@ 4.4. Ablation Study
The training of DreamPBR consists of some alternative modules and additional loss functions. In this section, we focus on evluating the effect of each of designs. To evaluate them, we randomly selected 100 textures from our obtained data that are not used in the whole training stage.
@@ 4.4.1. PBR Decoder
When PBR Decoder is trained, we introduce $\mathcal{L}_{\text{render}}$ to solve the regression problem from images rendered with random lights and viewpoints, which enforces the decoded textures is realistic after being rendered. It reduces the search space of output values compared to the one that rendering images is not used. We trained two PBR decoder with and without $\mathcal{L}_{\text{render}}$, and evaluate the effectiveness of them by comparing the outputs with reference textures. Figure 14 presents the comparsion results, in which our rendering-aware decoder is capable of achieving more realistic results in rendered results and more consistent reuslts in generated textures.
@@ 4.4.2. Super-Resolution Module
Although the initially super-resolution models shows greate reuslts in natural images, we finetune it again with our material data and employ a noval rendering loss $\mathcal{L}_{\text{render}}$ from the level of perception. In practice, we finetune super-resolution modules for each component of textures based on the pre-trained Real-ESRGAN as our baseline. With four single modules(albedo, metallic, normal, and roughness), we jointly finetune them and introduce the $\mathcal{L}_{\text{render}}$ by rendering four textures after super resolution to image space. The comparison results are shown in Figure 15. Similar to the training of PBR Decoder, the finetuning super-resolution modules with $\mathcal{L}_{\text{render}}$ contributes to better results.
@@ 4.4.3. Highlight-aware decoder
As mentioned in Section 3.3, we introduce a highlight-aware albedo decoder to remove the potential highlights in generated RGB images. For a good de-highlight module, there are two keypoints to be taken into account: 1) effectively removing the highlights in images, and 2) leaving unchanged for those without highlights. In practice, only training on rendered image potentially affect the decoded albedo(without highlights), so that we finetune the highlight-aware deocder by randomly choosing rendered images from different lights or pure albedo maps. Furthermore, we compare the outputs of highlight-aware decoder with the ones of initially pre-trained decoder in Figure 16, suggesting that our decoder addresses the issues of two keypoints above.
@@ 4.4.4. Pixel Control
To reliaze the sketche-guidance control, we embed a pre-trained ControlNet in DreamPBR. However, different with IP-Adapter for Style Control that focus on incorporating semantics of images in clipspace independent of training data, the initially ControlNet leads to domain shift, from albedo domain back to image domain, in our experiments. To address this problem, we finetune the ControlNet with our sketch-albedo pairs as mentioned above. The comparison of ControlNet before and after being finetuned is shown in Figure 17.
@@ 4.5. Limitations
Despite the promising capabilities of DreamPBR in generating high-quality and diverse material textures, our method encounters certain limitations that merit further exploration and improvement. We employ normal maps to reveal surface details in textures. However, using of normal maps without displacement maps leaves self-occlusion ignored when rendering them with those textures, which makes the rendering results unrealistic. In addtion, although a more lengthy description contributes to a more detail texture that user want, it is also a complex work for users to produce such a detail description like “a PBR material of wall, concrete wall, outdoor, cracked, man made, rough, painted…”.
@@ 5. Conclusions and Future work
In this paper, we propose DreamPBR, a novel diffusion-based generative framework for creating physically-based material textures. Our methods do not reply on large data sets as image generation does but transfer its original prior information to desired textures. Given text descriptions and other optional multi-modal conditions, we can generate textures that are highly consistent with text descriptions and the other condtions such as styles of RGB images and patterns of binary images. By using DreamPBR, one can create planar textures freely according to their imaginations. Specifically, we start with finetune a diffusion models for albedo generation and then decompose albedo to other SVBRDFs(normal, metallic, and roughness) by our highlight-aware decoder and PBR Decoder. For higher resolution texutres, we easily introduce a additional loss function in rendering image to our super-resolution module and bring significant improvement visually. With the properties above, DreamPBR can also produce some textures for simple geometries by dialogue with LLM.
For the future work, although DreamPBR currently targets planar textures, it’s could be extended to complex geometries with further development of retopology. Additionally, because of our effective PBR Decoder and highlight-aware deocder, DreamPBR is potentially to be used in SVBRDF esitmation. Lastly, there are inevitably problems such as limited resolution and time-consuming inference when utilizing diffusion models, which is also a challenging problem in the future.

### v2

@@ 4.2. Generation Results
DreamPBR is capable of generating realistic or magic materials with only descriptions. To demonstrate the ability to synthesize wide materials, we obtain a mount of descriptions of materials in by LLM for each type, which is used to sample materials with DreamPBR. The generated textures are enhanced by the super-resolution module and are then rendered as shown in Figure 3. In our sampled 400 textures, they show high consistency with text and the mean of CLIP Score between rendering images and given prompts is 30.198. Besides the consistency of text and images, the diversity of results is quite important for text-driven generative models as well. As demonstrated in Figure 7, we further sample several textures with the same prompt but different random seeds, DreamPBR succeeds in producing diverse textures that follow the descriptions we specify.
@@ 4.2.1. Tileable texture generation
Although the users would introduce various controls, we can generate seamless tileable textures all the time, which allows users to apply the generated textures in different scales and different scenes. In Figure 8, we present several tileable textures from direct and guided generation with their splicing results, showing the effectiveness of circular padding in our method as mentioned in Section 3.2.
@@ 4.2.2. Results of Pixel Control
By finetuning an additional ControlNet, DreamPBR is able to generate textures according to given patterns. In practice, a designer could decide on a pattern in advance, and then try different materials. It may also be the other way around. For those two situations, DreamPBR ensures reasonable textures for certain patterns or materials as demonstrated in Figure 4 and Figure 5.
With additional control of binary images, inpainting is also a usual method for users to obtain specified results so we present several inpainting results in Figure 9 to replace a region in texture with another region users describe.
@@ 4.2.3. Results of Style Control
A styled image expresses more easily for a person than only text like Su et al. (2021) does. To do so, we evaluate the adaptation of Su et al. (2021) for our Style Control. Specifically, we obtain several styled images online and present the generation results under different styles from images as shown in Figure 6. Figure 10 illustrates the situation in that users would like to combine Style Control with Pixel Control, which enables users to generate the results they want more freely.
@@ 4.2.4. Results of Shape Control
With the ability to generate various textures, DreamPBR can be extended to non-planar objects such as chairs. Specifically by giving a segmented object, we can utilize dialogue with a large language model to get different descriptions of each region. For more specified objects, a more direct way is to be in conjunction with cropped areas from exemplar images used with pixel control and style control. Thanks to the tileable features, the results from our pipeline of Shape Control are shown in Figure 11.
@@ 4.3. Comparative Experiments
Leveraging the state-of-the-art generative model, StableDiffusion, DreamPBR is very competitive with previous methods for materials generation. We compare the results generated from DreamPBR of different materials against MaterialGAN  (Guo et al., 2020) and TileGen  (Zhou et al., 2022) in Figure 12.  Notably, there are only two categories provided in the competing methods so our results are generated by giving prompts, “a PBR material of ground, stone” and “a PBR material of metal”. The comparison shows that DreamPBR can generate textures following the distribution of realistic data from datasets like GAN-based methods as well as magic textures from prior information for 2D images.
Moreover, we compare our Pixel Control with those of TileGen in generation with sketches guidance. The comparison results are shown in Figure 13, in which we demonstrate different generation results of TileGen and ours with the same binary masks. DreamPBR surpasses TileGen in sketches-driven generation and shows fewer artifacts and more precise controls than previous research on material generation like TileGen.
@@ 4.4. Ablation Study
The training of DreamPBR consists of some alternative modules and additional loss functions. In this section, we focus on evaluating the effect of each of the designs. To evaluate them, we randomly selected 100 textures from our obtained data that were not used in the whole training stage.
@@ 4.4.1. PBR Decoder
When the PBR Decoder is trained, we introduce $\mathcal{L}_{\text{render}}$ to solve the regression problem from images rendered with random lights and viewpoints, which enforces that the decoded textures are realistic after being rendered. It reduces the search space of output values compared to the one that rendering images is not used. We trained two PBR decoders with and without $\mathcal{L}_{\text{render}}$, and evaluated their effectiveness of them by comparing the outputs with reference textures. Figure 14 presents the comparison results, in which our rendering-aware decoder is capable of achieving more realistic results in rendered results and more consistent results in generated textures.
@@ 4.4.2. Super-Resolution Module
Although the super-resolution models originally show great results in natural images, we finetune it again with our material data and employ a novel rendering loss $\mathcal{L}_{\text{render}}$ from the level of perception. In practice, we finetune super-resolution modules for each component of textures based on the pre-trained Real-ESRGAN as our baseline. With four single modules(albedo, metallic, normal, and roughness), we jointly finetune them and introduce the $\mathcal{L}_{\text{render}}$ by rendering four textures after super-resolution to image space. The comparison results are shown in Figure 15. Similar to the training of PBR Decoder, the finetuning super-resolution modules with $\mathcal{L}_{\text{render}}$ contributes to better results.
@@ 4.4.3. Highlight-aware decoder
As mentioned in Section 3.3, we introduce a highlight-aware albedo decoder to remove the potential highlights in generated RGB images. For a good de-highlight module, there are two key points to be taken into account: 1) effectively removing the highlights in images, and 2) leaving them unchanged for those without highlights. In practice, only training on rendered images potentially affects the decoded albedo(without highlights), so we finetune the highlight-aware decoder by randomly choosing rendered images from different lights or pure albedo maps. Furthermore, we compare the outputs of the highlight-aware decoder with the ones of the initially pre-trained decoder in Figure 16, suggesting that our decoder addresses the issues of two key points above.
@@ 4.4.4. Pixel Control
To realize the sketch-guidance control, we embed a pre-trained ControlNet in DreamPBR. However, different from the IP-Adapter for Style Control focuses on incorporating semantics of images in clip space independent of training data, the initial ControlNet leads to domain shift, from the albedo domain back to the image domain, in our experiments. To address this problem, we finetuned the ControlNet with our sketch-albedo pairs as mentioned above. The comparison of ControlNet before and after being finetuned is shown in Figure 17.
@@ 4.5. Limitations
Despite the promising capabilities of DreamPBR in generating high-quality and diverse material textures, our method encounters certain limitations that merit further exploration and improvement. We employ normal maps to reveal surface details in textures. However, using normal maps without displacement maps leaves self-occlusion ignored when rendering them with those textures, which makes the rendering results unrealistic. In addition, although a more lengthy description contributes to a more detailed texture that the user wants, it is also complex work for users to produce such a detailed description like “a PBR material of the wall, concrete wall, outdoor, cracked, man-made, rough, painted…”.
@@ 5. Conclusions and Future work
In this paper, we propose DreamPBR, a novel diffusion-based generative framework for creating physically-based material textures. Our methods do not rely on large data sets as image generation does but transfer their original prior information to desired textures. Given text descriptions and other optional multi-modal conditions, we can generate textures that are highly consistent with text descriptions and the other conditions such as styles of RGB images and patterns of binary images. By using DreamPBR, one can create planar textures freely according to their imagination. Specifically, we start with finetuning diffusion models for albedo generation and then decompose albedo to other SVBRDFs(normal, metallic, and roughness) by our highlight-aware decoder and PBR Decoder. For higher-resolution textures, we easily introduce an additional loss function in rendering images to our super-resolution module and bring significant improvement visually. With the properties above, DreamPBR can also produce some textures for simple geometries by dialogue with LLM.
For future work, although DreamPBR currently targets planar textures, it could be extended to complex geometries with further development of retopology. Additionally, because of our effective PBR Decoder and highlight-aware decoder, DreamPBR has the potential to be used in SVBRDF estimation. Lastly, there are inevitably problems such as limited resolution and time-consuming inference when utilizing diffusion models, which is also a challenging problem in the future.